A rectangle has an area of 24 square units. The width is 5 units less than the length. What is the length, in units, of the rectangle?
Answer by RAY100(1637)

Area = Length * Width
.
but W = L -5
.
A= L*W = L*(L-5) = L^2 -5L
.
but we are given A= 24,,,,so,,24 = L^2 -5L
.
or L^2 -5L -24 =0,,,,,factoring
.
(L-8)(L+3) =0
.
L=8,,,-3(not realsitic)
.
W= L-5 = 8-5 = 3
.
Check: A=L*W= 8*3 = 24,,,,,ok